Electric fields up to 2.00 ✕ 105 N/C have been measured inside of clouds during electrical storms. Neglect the drag force due to the air in the cloud.

(a) What acceleration does the maximum electric field produce for protons in the cloud? Express your answer in SI units.
  

Express your answer as a fraction of g.


(b) If the electric field remains constant, how far will the proton have to travel to reach 5% of the speed of light if it started with negligible speed?

(c) Can you neglect the effects of gravity? Explain your answer
.
Answer #1

A.

As

F = q E

and

a = F / m

Then

a = q E / m

Thus,

a = 1.9185*10^13 m/s^2 or 1.96*10^12 g [ANSWER]

********************

B.

If we treat this classically,

v = vo + at

--> t = v / a when vo = 0.

Thus,

t = 0.05 c / a = 7.813*10^-7 s [ANSWER]

***********************

C.

Yes, because the acceleration is much greater than g.